Monsoon Inc. is considering bidding on a government project. To do the project, the company must make an initial investment of $8 million to purchase the necessary equipment. The project will last for five years, at the end of which the equipment can be salvaged for $500,000. The equipment has a CCA rate of 30%. The bidding process for the project requires the firm to submit a bid for a constant amount of $X before-tax, to be remitted by the government to the winning bidder each year. The firm’s marginal tax rate is 40%, and the required rate of return on similar projects is 18%. What is the minimum bid that the firm should submit for this project? Round your answer to the nearest dollar.

Select one:

a. $8,000,000
b. $3,191,717
c. $1,915,030
d. $3,308,198
e. $5,988,626
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

ANSWER : b : $3,191,717

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum , 11.A. A.--- 9- ー E ゴText ずWrap Text General ! Fill. ey Format Painter B 1 U. ㄧ˙ タ.Δ. -=- i 便困Merge & Center. $,%, ,,oan conditional Format Cell Insert Delete Format Insert Delete Format 2 ClearFe Select Formatting ▼ as Table, Styles ▼ Clipboard Font Alignment Number Cells Edting 038 IJ IK IM IN lO IP Ia Is IT IU 38 39 STEP 1 40 41 42 43 CALCULATION OF PRESENT VALUE OF TAX SHIELD PVCCATS = 45 46 47 48 49 50 51 52 53 54 C COST OF EQUIPMENT d DEPRECIATION RATE Tc = TAX RATE S- SALVAGE VALUE r DISCOUNT RATE 8000000 30% 40% 500000 18% PVCCATS ((( 8000000*0.3*0.4)/(0.18+0.3) ((1+0.5*0.18)/(1+0.18)) - (500000*0.3*0.4)/(0.3 +0.18) (1/((1+0.18)A5) VCCATS = 1792818.98 56 HPR GMAM EAC MACRS LEASE RATIOCASHBUDGETwacc BOND EPS EBIT REPLACEMENT STats NP BETA LEVERED 29-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Font Alignment Number Cells 075 IJ IK IM IN lO IP IR Is IT IU 57 58 STEP 2 PRESENT VALUE OF ALL CASH FLOWS 59 60 61 62 STEP 3 INTIAL CASH OUTFLOW 63 64 65 66 STEP 4 NPV STEP 1+ STEP 2 -STEP 3 67 68 69 70 71 72 73 74 75 CF( 1-0.4) X PVI FA @ 18%, 5 YEARS + (SALVAGE VALUEJX PVIF @ 18%, 5 YEARS) CF*(1-0.4) X 3.12717102 (500000) X 0.43710922 (CFX 1.87630261) +218554.61 COST OF MACHINE NPV = 1792818.98 + CF X 1.87630261 + 218554.61-8000000 0 1792818.98 + CFX 1.87630261 218554.61 -8000000 0= CFX 1.87630261-5988626.41 CF 3191717 NOW WE PUT NPV 0 AND SOLVE IT REQUIRED ANSWER HPR GMAM EAC MACRS RATİOCASHBUDGET i Wacc BOND s EBIT . REPLACEMENT , STats . NPV BETA LEVERED . 02:50